OpenStudy (anonymous):

The ages if Sylvana's three daughters are consecutive even integers. Find their ages if the sum of their ages is 60.

OpenStudy (anonymous):

@Luigi0210

OpenStudy (texaschic101):

3 consecutive even integers.. x x + 2 x + 4 the sum of their ages is 60 x + (x + 2) + (x + 4) = 60 -- combine like terms 3x + 6 = 60 -- subtract 6 from both sides 3x = 54 -- divide both sides by 3 x = 18 x + 2 = 18 + 2 = 20 x + 4 = 18 + 4 = 22 check.. 18 + 20 + 22 = 60 60 = 60 (correct) so their ages are : 18,20,and 22

OpenStudy (arivillafana):

age of first daughter ->x age of second daughter ->x+2 age of third daughter ->x+4 x+x+2+x+4=60 3x+6=60

OpenStudy (triciaal):

consecutive means next in order even count by 2's

OpenStudy (triciaal):